HW 1c Question 2 (of 8) value: 10.00 points 1 out of 3 attempts Enter your answer in the provided box. Assistance The normal lead content in human blood is about 0,40 part per million (that is, 0.40 g of lead per million grams of blood). A value of 0.80 part per million (ppm) is considered to be dangerous. How many grams of lead are contained in 6.0 x 10 g of blood (the amount in an average adult) if the lead content is 0.6 ppm? Check My Work View Question Show Me Guided Solution Question Help Report a Proble 渋: 3g
Answer #1

Given that 0.4 ppm of Lead in Blood = 0.4 g of Lead in 1000000 g of Blood

So, 0.62 ppm of Lead in Blood = 0.62 g of Lead in 1000000 g of Blood

Conversely, 1000000 g of Blood = 0.62 g of Lead

Implies, 1 g of Blood = (0.62/1000000) g of Lead

Therefore, 6000 g of Blood = [(0.62 \times 6000)/1000000] g of Lead

                                      = 3.72 \times 10-3 g of Lead is present in 6 \times 103 g of Blood